# Reinforcement Learning with Intrinsically Motivated Feedback Graph for Lost-sales Inventory Control

### v1

@@
Reinforcement learning (RL) has proven to be well-performed and general-purpose in the inventory control (IC). However, further improvement of RL algorithms in the IC domain is impeded due to two limitations of online experience. First, online experience is expensive to acquire in real-world applications. With the low sample efficiency nature of RL algorithms, it would take extensive time to train the RL policy to convergence. Second, online experience may not reflect the true demand due to the lost sales phenomenon typical in IC, which makes the learning process more challenging. To address the above challenges, we propose a decision framework that combines reinforcement learning with feedback graph (RLFG) and intrinsically motivated exploration (IME) to boost sample efficiency. In particular, we first take advantage of the inherent properties of lost-sales IC problems and design the feedback graph (FG) specially for lost-sales IC problems to generate abundant side experiences aid RL updates. Then we conduct a rigorous theoretical analysis of how the designed FG reduces the sample complexity of RL methods. Based on the theoretical insights, we design an intrinsic reward to direct the RL agent to explore to the state-action space with more side experiences, further exploiting FG’s power. Experimental results demonstrate that our method greatly improves the sample efficiency of applying RL in IC. Our code is available at https://anonymous.4open.science/r/RLIMFG4IC-811D/
@@ 1 Introduction
Inventory control (IC) is a crucial and practical problem, serving as the basis of business efficiency for supply chain management. IC problems are challenging due to the intractability caused by the inherent complexity and the difficulty of finding optimal solutions within a reasonable timeframe. In the past few decades, researchers Zipkin (2008a), Bai et al. (2023), Xin (2021), and Chen et al. (2024) have designed a few heuristic methods based on specific model assumptions. However, such methods often face challenges due to the curse of dimensionality (Goldberg et al. 2021), where the problem size grows exponentially as the lead time increases. Here lead time refers to the duration between placing and receiving an order. Another limitation is that these model-based methods are not flexible enough to generalize to various environmental settings. These limitations call for a more adaptable approach to IC problems, for which data-driven control methods stand out as a promising alternative.
@@
Reinforcement learning (RL) has gained significant attention as a powerful data-driven technique for solving complex sequential decision-making problems. In particular, it offers several advantages for addressing the challenges of IC problems. Firstly, RL allows for the discovery of optimal policies without relying on strong problem-specific assumptions, enabling more generalizable solutions (Nian et al. 2020). Secondly, when combined with the deep neural network, Deep RL (DRL) can handle large state and action spaces, making it suitable for problems with high-dimensional state variables (Dehaybe et al. 2024). Early works demonstrate the feasibility of RL in IC problems. Oroojlooyjadid et al. (2022) showcases the ability of a Deep Q-network (DQN) to discover near-optimal solutions for the widely recognized beer distribution game. Gijsbrechts et al. (2022) introduces the A3C algorithm into IC problems and show that it can achieve acceptable performance but not better than heuristic methods. Stranieri and Stella (2023) benchmarks various DRL methods such as A3C, PPO, and vanilla policy gradient (VPG) in IC problems. More recently, researchers tend to explore how DRL can address various scenarios within the IC problems. These scenarios include: non-stationary uncertain demand (Dehaybe et al. 2024, Park et al. 2023), multi-product (Sultana et al. 2020, Selukar et al. 2022), variable kinds of products (Meisheri et al. 2020, 2022), multi-echelon supply chains (Wu et al. 2023, Stranieri et al. 2024), one-warehouse multi-retailer (Kaynov et al. 2024), and the stochastic capacitated lot sizing problem (Van Hezewijk et al. 2023).
@@
However, traditional RL methods are characterized by low sample efficiency, which becomes a significant barrier when implementing these techniques in real-world IC scenarios, as obtaining experiences can be both costly and time-consuming (Boute et al. 2022, De Moor et al. 2022). Furthermore, this sample inefficiency issue is enlarged in lost-sales IC problems because of censored demands (Chen et al. 2024), which refers to the phenomenon when customers’ real demands are unobservable due to insufficient inventory. For instance, if the order is placed daily, it takes over a year to generate four hundred experiences and part of them may be censored, making them too few to update the RL policy. De Moor et al. (2022) tries to alleviate this problem by incorporating heuristic knowledge, such as the base-stock policy, into DQN with reward-shaping. Although this method can improve the sample efficiency, it still relies on specific model heuristics, making it hard to generalize to different scenarios. Overall, resolving the low sample efficiency of RL without strong heuristics is crucial in solving IC (especially lost-sales IC) problems. The detailed related work is in Appendix 6.
@@
By addressing the above-mentioned limitation, this paper proposes a novel decision framework that combines the reinforcement learning with feedback graphs (RLFG) and intrinsically motivated exploration (IME):
@@
1) We tailor the feedback graph (FG) based the general property of lost-sales IC problems rather than strong heuristics (e.g. known demand distribution). In particular, the connectivity of FG is adjusted dynamically based on the relationship between the demand and inventory rather than being static in the environment. With FG, the sample efficiency in the training process is significantly improved by allowing the agent to acquire not only online experiences but also side experiences from FG.
@@
2) We conduct a theoretical analysis of how FG reduces the sample complexity with Q-learning as an example. It demonstrates that FG decreases the sample complexity by improving the update probabilities across all state-action pairs.
@@
3) Inspired by these theoretical insights, we design a novel intrinsic reward that guides the RL algorithm to explore towards the state-action space where more side experiences can be obtained thereby further boosting sample efficiency.
4) We evaluate our method on the standard discrete lost-sales inventory control environment. Our empirical results demonstrate the superior sample efficiency improvement by FG and the intrinsic reward separately, underlining the effectiveness of our design.
2 Background and Problem Formulation
2.1 Reinforcement Learning with Feedback Graph
RLFG is proposed by Dann et al. (2020) to reduce the sample complexity of RL. In typical RL scenarios, an agent can only get one experience each time as feedback that can be used to update the policy. However, when some prior knowledge about the environments is available, it becomes possible to observe additional experiences involving other states and actions. RLFG aims to explore how RL algorithms can benefit from these side experiences by constructing a feedback graph (FG). Here “side experiences” bears the same meaning as “side observations” in Dann et al. (2020), and we use the term “side experiences” to avoid confusion with “observations” in partially observed MDPs. FG is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set $\mathcal{V}=\{\bm{v}|\bm{v}=(\bm{s},a)\}$ and $\mathcal{E}$ is the edge set =$\mathcal{E}=\{\bm{v}\rightarrow\bm{\bar{v}}\}$, formalized by the side information indicating that if the agent visits $\bm{v}$, it can also observe other vertices $\bm{\bar{v}}$. The total experiences $\mathcal{O}_{t}$ observed by the agent at $t$ from $\mathcal{G}$ is
$\mathcal{O}_{t}(\mathcal{G})=\{(\bm{s}_{t},a_{t},r_{t},\bm{s}_{t+1})\}\cup\{\bm{\bar{s}_{t}},\bar{a_{t}},\bar{r_{t}},\bm{\bar{s}}_{t+1}\}.$ (1)
@@ 2.2 Lost Sales Inventory Control Problem
Table 1: The demands and state variables.
Symbol Representation Symbol Representation
$L\geq 0$ Order lead time $a_{t}\in\mathbb{N}$ Order at time $t$
$t=1,...,T$ Time index $a^{\text{max}}\in\mathbb{N}$ Maximum order
$d_{t}\in\mathbb{N}$ Real demand in period $t$ $c\in\mathbb{R}$ Unit cost of procurement
$d^{o}_{t}\in\mathbb{N}$ Demand observed in period $t$ $h\in\mathbb{R}$ Unit cost of holding inventory
$d^{\text{max}}\in\mathbb{N}$ Maximum demand $p\in\mathbb{R}$ Unit cost penalty of lost sales
$\bm{s}_{t}\in\mathbb{N}^{L}$ State: $(y_{t},a_{t+1-L},...,a_{t-1})$ $\gamma\in(0,1]$ Discount factor
$y_{t}\in\mathbb{N}$ Inventory at time $t$, $y^{\text{max}}\in\mathbb{N}$ Maximum inventory at time $t$,
after the order due at $t$ arrives after the order due at $t$ arrives
We formulate a standard, single-item, discrete-time, and lost-sales IC problem following Zipkin (2008a), Gijsbrechts et al. (2022) and Xin (2021) with environment variables in Table 1. The “lost-sales” means that customers leave if there is not enough inventory without any way to record the excess demand. The lost-sales IC problem considers three kinds of costs in each time step $t$, which are the cost of procurement $f_{1}(\bm{s}_{t})$, the cost of holding inventory $f_{2}(\bm{s}_{t})$, and the cost penalty of lost sales $f_{3}(\bm{s}_{t})$. The cost terms are specified below, where $[x]^{+}=\max(x,0)$.
$f_{1}(\bm{s}_{t},a_{t})=ca_{t},f_{2}(\bm{s}_{t})=h[y_{t}-d_{t}]^{+},f_{3}(\bm{s}_{t})=p[d_{t}-y_{t}]^{+}.$ (2)
@@
Here, the term “censored” indicates the otherwise case in Equation 3, when we can only observe all inventory is sold but do not know the value of $d_{t}$. The objective is to minimize the total cost over the time horizon $T$ under the uncertainty on the demand side, which is
$\min_{\{a_{t}|t=0,...,T\}}\sum_{t=0}^{T}\gamma^{T-t}f(\bm{s}_{t},a_{t})=\sum_{t=0}^{T}\gamma^{T-t}[f_{1}(\bm{s}_{t},a_{t})+f_{2}(\bm{s}_{t})+f_{3}(\bm{s}_{t})].$ (4)
@@ 3 Method
Figure 1: IC decision framework based on RLFG and IME. Solid lines depict sampling steps and dashed lines represent model-updating steps.
3.1 IC Decision Framework with RLFG and IME
The proposed decision framework for IC problems aims to enhance the sample efficiency of off-policy RL algorithms by integrating RLFG and IME together. This framework relies on two key assumptions. First, the real demand is unknown and the observed demand may be censored. Second, the experiences generated in real-world operations are limited due to the cost involved in collecting online data. As illustrated in Figure 1, the FG module generates side experiences to improve the sample efficiency and the intrinsic reward module aids the exploration to further exploit the power of the FG module. Here, the RL module can be any off-policy RL algorithm, such as DQN, DDPG, Rainbow, or TD3. Note that experiences from both reply buffers are sampled with no differences in the model-updating step. Then a new set of side experiences are generated based on the sampled experiences. These new side experiences are just used to calculate intrinsic rewards of the sampled experiences. Here we provide Rainbow-FG as an example in Algorithm 1.
Algorithm 1 Rainbow-FG Algorithm.
1:  Initialize the distributional Q-network $Q$ with parameter $\theta$, the target Q-network $Q^{\prime}$ with parameter $\phi$, the replay buffer $\mathcal{R}$, and the side replay buffer $\mathcal{R}^{s}$
2:  for episode = $1,...,M$ do
3:     Initialize the environment state $s_{0}$
4:     for $t=0,...,T$ do
5:        Get the order $a_{t}=\begin{cases}\text{argmax}_{a}\ Q(\bm{s}_{t},a),&\ w.p\ \ 1-\epsilon\\ \text{a random action},&\ w.p\ \ \epsilon\end{cases}$
6:        Execute action $a_{t}$ and get next state $\bm{s}_{t+1}$ and reward $r_{t}$
7:        Get the observed demand $d_{t}^{o}$
8:        Get the side experiences $\{(\hat{\bm{s}_{t}},\hat{a_{t}},\hat{r_{t}},\hat{s}_{t+1},d^{o}_{t})\}\to$ Algorithm 2
9:        $\mathcal{R}=\mathcal{R}\cup(\bm{s}_{t},a_{t},r_{t},\bm{s}_{t+1},d_{t}^{o})$ and $\mathcal{R}^{s}=\mathcal{R}^{s}\cup\{(\hat{\bm{s}_{t}},\hat{a_{t}},\hat{r_{t}},\hat{\bm{s}}_{t+1},d_{t}^{o})\}$
10:        Model Update:
11:        Randomly sample a batch of experiences $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}$ from $\mathcal{R}\cup\mathcal{R}^{s}$
12:        Get all side experiences $\{(\hat{\bm{s}},\hat{a},\hat{r},\hat{\bm{s}}^{\prime},d^{o})\}$ for each experience in $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}\to$ Algorithm 2
13:        Get the intrinsic reward $r^{in}$ of this batch of experiences $\to$ Algorithm 3
14:        Get the final reward $r=(1-\beta)\times r+\beta\times r^{in}$
15:        Update parameters
@@ 3.2 Feedback Graph in Inventory Control
Motivated by the reduction in sample complexity by FG (Dann et al. 2020), we incorporate FG into the IC problem. FG is naturally suitable for the IC problem since the IC problem is a structured MDP with most environmental transition components determined and predictable depending on the demand. Furthermore, the demand is usually independent of the state-action space. If we can observe the demand, the experiences of other state-action pairs can also be obtained through leveraging the underlying properties of such structured MDPs. Thus the main challenge is how to construct FG considering the lost-sales property, whose demand is potentially censored. To solve this problem, we propose to construct FG dynamically based on the observed demand. If the observed demand is real, FG generates side experiences of all other state-action pairs. Even if the observed demand is censored, FG can generate side experiences of the state-action pairs whose inventory is less than the observed demand.
@@
Algorithm 2 shows the details of the FG module. When $d^{o}_{t}$ is uncensored, FG is a complete graph and all state-action pairs can be used to generate side experiences based on $d^{o}_{t}$. When $d^{o}_{t}$ is censored, it becomes more complex. For the state-action pairs having larger inventory numbers than that in $\bm{s}_{t}$, $d^{o}_{t}$ is not a correct demand for them and unfortunately $d_{t}$ is unknown. Still using $d^{o}_{t}$ will generate wrong side experiences. We can only use the state-action pairs having smaller inventory numbers than that in $\bm{s}_{t}$ to generate side experiences, which means that FG is a partially connected graph. Note that FG is still dynamically constructed based on $d^{o}_{t}$ under the censored case since different $d^{o}_{t}$ generates different numbers of side experiences.

### v2

@@
Reinforcement learning (RL) has proven to be well-performed and versatile in inventory control (IC). However, further improvement of RL algorithms in the IC domain is impeded by two limitations of online experience. First, online experience is expensive to acquire in real-world applications. With the low sample efficiency nature of RL algorithms, it would take extensive time to collect enough data and train the RL policy to convergence. Second, online experience may not reflect the true demand due to the lost-sales phenomenon typical in IC, which makes the learning process more challenging. To address the above challenges, we propose a training framework that combines reinforcement learning with feedback graph (RLFG) and intrinsically motivated exploration (IME) to boost sample efficiency. In particular, we first leverage the MDP structure inherent in lost-sales IC problems and design the feedback graph (FG) tailored to lost-sales IC problems to generate abundant side experiences aiding in RL updates. Then we conduct a rigorous theoretical analysis of how the designed FG reduces the sample complexity of RL methods. Guided by these insights, we design an intrinsic reward to direct the RL agent to explore to the state-action space with more side experiences, further exploiting FG’s capability. Experimental results on single-item, multi-item, and multi-echelon environments demonstrate that our method greatly improves the sample efficiency of applying RL in IC. Our code is available at https://github.com/Ziffer-byakuya/RLIMFG4IC
@@ 1 Introduction
Inventory control (IC) is a practical problem rooted in the real-world application of supply chain management (SCM). It mainly aims at minimizing operational costs while satisfying the demand of customers. The inherent complexity of modeling supply chain dynamics and demand uncertainty makes finding the optimal solutions for IC computationally intractable within reasonable time frames. To address this challenge, researchers (Zipkin 2008a, Bai et al. 2023, Xin 2021, Chen et al. 2024) have designed heuristic methods based on specific model assumptions over the past few decades. However, these methods face limitations in practice. Firstly, they struggle with the curse of dimensionality (Goldberg et al. 2021). As the lead time from order placement to reception increases, the problem size grows exponentially. Secondly, their reliance on specific assumptions may restrain their adaptability; they cannot easily generalize across diverse environmental settings. For example, most well-performed heuristic methods for single-item IC cannot be extended to multi-item or multi-echelon IC. These limitations call for a more adaptable approach to IC problems, for which data-driven control methods stand out as a promising alternative.
@@
Reinforcement learning (RL) has shown promise in handling complex sequential decision-making problems. In particular, it offers several advantages for addressing the challenges of IC problems. Firstly, RL allows for the discovery of optimal policies without relying on strong problem-specific assumptions, enabling more generalizable solutions (Nian et al. 2020). Secondly, when combined with the deep neural network, Deep RL (DRL) can handle the high-dimensional state and action spaces typical of IC, thereby alleviating the curse of dimensionality (Dehaybe et al. 2024). Early works have demonstrated the feasibility of RL in IC problems. For instance, Oroojlooyjadid et al. (2022) showcases the ability of Deep Q-network (DQN) to discover near-optimal solutions for the widely recognized beer distribution game (a type of game for SCM simulation that illustrates the complexities and challenges of IC). Gijsbrechts et al. (2022) achieves performance comparable to that of heuristic methods using the A3C algorithm in IC problems. Stranieri and Stella (2023) benchmarks various DRL methods such as A3C, PPO, and vanilla policy gradient (VPG) in IC problems. More recently, researchers have explored DRL for various IC scenarios, including non-stationary uncertain demand (Dehaybe et al. 2024, Park et al. 2023), multi-product (Sultana et al. 2020, Selukar et al. 2022), variable kinds of products (Meisheri et al. 2020, 2022), multi-echelon supply chains (Wu et al. 2023, Stranieri et al. 2024), one-warehouse multi-retailer (Kaynov et al. 2024), and the stochastic capacitated lot sizing problem (Van Hezewijk et al. 2023).
@@
However, traditional RL methods suffer from low sample efficiency, presenting a major barrier to real-world implementation where experience collection can be costly and time-consuming (Boute et al. 2022, De Moor et al. 2022). Furthermore, this issue is escalated in lost-sales IC problems because of censored demands (Chen et al. 2024), where customers’ real demands are unobservable due to insufficient inventory. For instance, if the order is placed daily, it takes over a year to collect four hundred experiences and part of them may be censored. The scarcity of valid experience presents a true challenge for training data-hungry RL policies. De Moor et al. (2022) tries to alleviate this problem by incorporating heuristic knowledge, such as the base-stock policy, into DQN with reward-shaping. Although this method can improve the sample efficiency, it still relies on specific model heuristics, making it hard to generalize to different scenarios. Therefore, overcoming the sample inefficiency of RL methods without assuming strong heuristics remains a key challenge for effectively applying these techniques to IC problems, especially under lost-sales consideration. The detailed related work is provided in Appendix 6.
@@
To address the above-mentioned limitations of DRL for IC problems, this paper proposes a novel training framework that combines reinforcement learning with feedback graphs (RLFG) and intrinsically motivated exploration (IME):
@@
1) We tailor the feedback graph (FG) based on the general property of lost-sales IC problems rather than case-specific heuristics (e.g. known demand distribution). In particular, the connectivity of FG is adjusted dynamically based on the relationship between the demand and inventory to be adaptive to the lost-sales property. With FG, the sample efficiency in the training process is significantly improved by allowing the RL agent to acquire not only online experiences but also side experiences from FG.
@@
2) We conduct theoretical analysis on how FG reduces the sample complexity with Q-learning to reveal its mechanism of improving the update probabilities across all state-action pairs in IC problems.
@@
3) Inspired by these theoretical insights, we design a novel intrinsic reward guiding the RL algorithm to explore towards the state-action space where more side experiences can be obtained thereby further boosting sample efficiency.
4) We evaluate our method across a wide range of IC settings, including the single-item, multi-item, and multi-echelon lost-sales inventory control environments. The empirical results demonstrate the superior sample efficiency and performance improvement of proposed methods, underlining the effectiveness of our designs and generalization ability to different scenarios.
2 Preliminary
@@ 2.1 Lost Sales Inventory Control Problem
Table 1: Nomenclature for environmental parameters.
Symbol Representation
$t=1,...,T$ Time index
$a_{t}\in\mathbb{N}$ Action: order at time $t$
$a^{\text{max}}\in\mathbb{N}$ Maximum order
$c\in\mathbb{R}$ Unit cost of procurement
$d_{t}\in\mathbb{N}$ Real demand at time $t$
$d^{o}_{t}\in\mathbb{N}$ Demand observed at time $t$
$d^{\text{max}}\in\mathbb{N}$ Maximum demand
$h\in\mathbb{R}$ Unit cost of holding inventory
$L\geq 0$ Order lead time
$p\in\mathbb{R}$ Unit cost penalty of lost sales
$\bm{s}_{t}\in\mathbb{N}^{L}$ State: $(y_{t},a_{t+1-L},...,a_{t-1})$
$y_{t}\in\mathbb{N}$ Inventory after receiving
the arrived order at time $t$.
$y^{\text{max}}\in\mathbb{N}$ Maximum inventory after receiving
the arrived order at time $t$.
$\gamma\in(0,1]$ Discount factor
We formulate a standard, single-item, discrete-time, and lost-sales IC problem following Zipkin (2008a), Gijsbrechts et al. (2022) and Xin (2021) with environment variables defined in Table 1. In lost-sales settings, excess demand is unrecorded when inventory is insufficient, with customers leaving due to unsatisfied demands. The lost-sales IC problem considers three kinds of costs in each time step $t$, which are the cost of procurement $f_{1}(\bm{s}_{t})$, the cost of holding inventory $f_{2}(\bm{s}_{t})$, and the cost penalty of lost sales $f_{3}(\bm{s}_{t})$. The cost terms are specified below,
$f_{1}(\bm{s}_{t},a_{t})=ca_{t},f_{2}(\bm{s}_{t})=h[y_{t}-d_{t}]^{+},$ (1)
$f_{3}(\bm{s}_{t})=p[d_{t}-y_{t}]^{+},\text{where }[x]^{+}=\max(x,0).$
@@
Here, the term “censored” refers to the case when $d_{t}>y_{t}$, i.e. we can only observe all inventory is sold but do not know the value of $d_{t}$. The objective is to minimize the total cost over the time horizon $T$ under the uncertainty on the demand side, which is
$\min_{\{a_{t}|t=0,...,T\}}$ $\sum_{t=0}^{T}\gamma^{t}f(\bm{s}_{t},a_{t})$ (3)
$=$ $\sum_{t=0}^{T}\gamma^{t}[f_{1}(\bm{s}_{t},a_{t})+f_{2}(\bm{s}_{t})+f_{3}(\bm{s}_{t})].$
@@ 2.2 MDP Formulation
2.3 Reinforcement Learning with Feedback Graph
RLFG is proposed by Dann et al. (2020) to reduce the sample complexity of RL. In typical RL scenarios, an agent can only get one experience each time as feedback that can be used to update the policy. However, given prior knowledge about the environments, additional side experiences involving other states and actions may be observable. RLFG explores how RL leverages these side experiences by constructing a feedback graph [^1] Here “side experiences” bears the same meaning as “side observations” in Dann et al. (2020), and we use the term “side experiences” to avoid confusion with “observations” in partially observable MDPs. (FG). FG is denoted as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set $\mathcal{V}=\{\bm{v}|\bm{v}=(\bm{s},a)\}$ and $\mathcal{E}$ is the edge set =$\mathcal{E}=\{\bm{v}\rightarrow\bm{\hat{v}}\}$, formalized by the side information indicating that if the agent visits experience $\bm{v}$, then the corresponding side experiences $\bm{\hat{v}}$ are also observable. The total experiences $\mathcal{O}_{t}$ observed by the agent at timestep $t$ from $\mathcal{G}$ is
$\mathcal{O}_{t}(\mathcal{G})=\{(\bm{s}_{t},a_{t},r_{t},\bm{s}_{t+1})\}\cup\{\bm{\hat{s}_{t}},\hat{a_{t}},\hat{r_{t}},\bm{\hat{s}}_{t+1}\}.$ (4)
@@ 3 Method
Figure 1: IC training framework based on RLFG and IME. Solid lines depict sampling steps and dashed lines represent model-updating steps.
3.1 IC Training Framework with RLFG and IME
We propose a training framework that incorporates RLFG and IME to enhance the sample efficiency of off-policy RL algorithms in IC problems. This framework relies on two primary assumptions causing sample scarcity. First, the real demand is unknown and may be censored. Second, the experiences collected in real-world operations are limited due to the cost of collecting online data. As illustrated in Figure 1, the FG module generates side experiences $\{(\mathbf{\hat{s}_{t}},\hat{a_{t}},\hat{r_{t}},\mathbf{\hat{s}_{t+1}},d^{o}_{t})\}$ to improve the sample efficiency and the intrinsic reward module aids the exploration by guiding the RL algorithm to explore in state-action space with more side experiences. Here, the RL module can be any off-policy RL algorithm, such as DQN, DDPG, Rainbow, or TD3. During model-updating steps, a batch of samples $\{(\mathbf{s},\hat{a},\hat{r},\mathbf{s^{\prime}},d^{o}_{t})\}$ are sampled from both reply buffers. Subsequently, the FG module generates a new set of side experiences $\{(\mathbf{\hat{s}},\hat{a},\hat{r},\mathbf{\hat{s}^{\prime}},d^{o}_{t})\}$, which are exclusively used to calculate the intrinsic rewards of the samples. Rainbow-FG is provided as an example in Algorithm 3 in Appendix 7.
@@ 3.2 Feedback Graph in Inventory Control
Motivated by the ability of FG to reduce sample complexity (Dann et al. 2020), we incorporate it into the IC problems. FG is naturally suitable since IC problems follow structured MDP formation and have predictable transitions based on demand, typically independent of state-actions pairs. Observing the demand allows us to leverage the underlying properties of structured MDPs to obtain side experiences. Thus the key question is how to construct FG considering the potentially censored demand due to lost sales. To address this problem, we propose to construct FG dynamically based on the observed demand.
@@
Algorithm 1 outlines the details of the FG module. When $d^{o}_{t}$ is uncensored, FG forms a complete graph allowing all state-action pairs to generate side experiences based on $d^{o}_{t}$. When $d^{o}_{t}$ is censored, only the state-action pairs with smaller inventory levels than $\bm{s}_{t}$ can be utilized to generate side experiences, resulting in a partially connected FG. Since $d^{o}_{t}$ is censored, the store can only observe that all the inventory are sold out and assume that less inventory can also be sold out without any information for more inventory. Note that FG is still dynamically constructed under the censored case as the number of side experiences depends on $d^{o}_{t}$.